# MF-GLaM: A multifidelity stochastic emulator using generalized lambda models

### v1

@@ 4.3 Multi-story building subject to an earthquake
In our last application, we aim to explore the applicability and performance of our framework in a real-world application involving real earthquake simulations, adopted from Zhu et al. (2023); Schär et al. (2025). More precisely, we consider a three-story steel frame structure subjected to stochastic seismic excitations. The structure is modeled as a three-degree-of-freedom system and is represented by a computational model, and the input is given by synthetic ground motion time series generated from a stochastic ground motion model fitted to a dataset of real ground motions. In this framework, the quantity of interest (QoI) is the maximum interstory drift, defined as the maximum of the individual interstory drifts, $\Delta_{1}(t),\Delta_{2}(t),\Delta_{3}(t)$, computed for each story during the seismic event:
@@
For the synthetic ground motion model, we adopt the model from Zhu et al. (2023), based on the models from Broccardo and Dabaghi (2017); Rezaeian and Der Kiureghian (2010). The uncertain seismic input is characterized by four ground motion parameters: the expected Arias intensity $I_{a}$, the time $t_{\text{mid}}$ at which 45% of the expected Arias intensity is reached, the effective duration $D_{5-95}$ of the ground motion, and the main frequency $\omega_{g}$ of the motion at $t_{\text{mid}}$. The ground motion parameters are statistically modeled using log-normal marginal distributions coupled with a Gaussian copula, with the details of the distributions shown in Table 6. As the four ground-motion parameters are high-level descriptors of an earthquake signal, they do not uniquely define a full time-series ground motion. Thus, a single parameter set can lead to significantly different possible signals.
@@
[…]
@@
In our study, different levels of fidelity are regulated by the time step of the simulations. The HF simulator uses a fine time step of $\delta t_{\text{H}}=0.01\leavevmode\nobreak\ \text{s}$, whereas the LF simulator is derived from the same model by increasing, but with a seven times larger time step $\delta t_{\text{L}}=0.07\leavevmode\nobreak\ \text{s}$. The average wall-clock runtime for an HF simulation is $\bar{t}_{\text{H}}=10.3\,\leavevmode\nobreak\ \text{s}$, whereas the average runtime for an LF simulation is $\bar{t}_{\text{L}}=1.8\,\text{s}$, meaning that an LF simulation is about 5.7 times faster than an HF simulation.
@@
As mentioned above, the four input parameters are statistically dependent and coupled via a Gaussian copula (see Table 6). However, for the PCEs used in constructing all GLaMs, the inputs are assumed to be independent, as detailed in Section 2. In such cases, it is common practice to disregard the input dependencies during model construction, as this often leads to more accurate pointwise predictions in data-driven settings compared to transforming the correlated inputs into uncorrelated ones via the highly nonlinear Rosenblatt transform (Torre et al., 2019). However, the input dependencies are still taken into account when generating training and validation samples.
Fig. 7 compares the empirical response distributions from our true HF and LF simulators to the corresponding predicted distributions from the HF-only GLaM (red curve) and the MF-GLaM (blue curve). Three random input samples obtained using LHS are shown in each row, and each column corresponds to a different HF training set size. For each input sample, we generated $N_{\text{R}}=500$ HF runs and $N_{\text{R}}=500$ LF runs to estimate the true HF (black dashed curve) and LF (gray dotted curve) distributions, respectively. The output has been multiplied by $1,\!000$ for convenient display in millimeters.
Despite the LF distribution often diverging from the corresponding HF response, the MF-GLaM successfully captures the bulk and shape of the HF distribution across various scenarios, especially when HF data are scarce. In contrast, the HF-only GLaM tends to be more sensitive to limited HF samples and can misalign with the empirical HF distribution, e.g., at $\boldsymbol{x}_{2}$ for $N_{\text{H}}=100$ or at $\boldsymbol{x}_{3}$ for $N_{\text{H}}=400$.
@@
Fig. 8 and the corresponding summary in Table 7 present the normalized Wasserstein distance $\epsilon_{W}$ for three surrogate models, LF-only, HF-only, and MF-GLaM, across increasing high-fidelity training set sizes from $N_{\text{H}}=100$ to $N_{\text{H}}=800$. The reference validation data were generated by running our HF simulator on 200 Latin hypercube samples, each with $N_{\text{R}}=250$ replications, resulting in a total of $50,\!000$ HF runs. In this application, the LF simulator is a numerically less refined version of the HF model, rather than a physically distinct one. This may explain why the LF-only GLaM performs relatively well at small HF training sizes ($N_{\text{H}}\leq 200$), when the available HF data are too limited to fully capture the underlying variability. Overall, the MF-GLaM (blue boxes) consistently achieves the lowest error among the three emulators, with especially large performance gains at smaller HF sample sizes ($N_{\text{H}}\leq 400$).
Accounting for the computational cost of HF and LF runs, where 5.7 LF runs are considered equivalent in cost to one HF run, the total cost of using $N_{H}=200$ HF samples and $N_{L}=1,\!000$ LF samples corresponds to $200+\frac{1,\!000}{5.7}\approx 375$ HF-equivalent runs. This configuration leads to an MF-GLaM with median $\epsilon_{\text{W}}=0.075$ (see Table 7). This error is significantly lower than the median $\epsilon_{\text{W}}$ of the HF-only GLaM trained on $N_{\text{H}}=400$ HF points, for which $\epsilon_{\text{W}}=0.094$, indicating that the MF approach performs better at a reduced total cost. A similar trend is observed for the MF-GLaM trained on $N_{H}=400$ and $N_{L}=1,\!000$ points, where the total cost amounts to $400+\frac{1,\!000}{5.7}\approx 575$ HF-equivalent runs. In this case, the MF-GLaM’s median error, $\epsilon_{\text{W}}=0.056$, is comparable to that of the HF-only GLaM trained on $N_{\text{H}}=800$ HF points, for which $\epsilon_{\text{W}}=0.059$. Thus, comparable performance is achieved at a significantly lower computational cost, corresponding to a budget saving of approximately $28\%$.
Figure 8: Multi-story building subject to an earthquake. Comparison of the normalized Wasserstein distance $\epsilon_{\text{W}}$ for LF-only, HF-only, and MF GLaMs across increasing HF experimental design sizes $N_{\text{H}}$. The box plots correspond to 25 repetitions of the procedure, and $\epsilon_{\text{W}}$ is computed using $200$ test points, with $250$ replications performed at each point.
Table 7: Multi-story building subject to an earthquake. Median (interquartile range) of the normalized Wasserstein distance $\epsilon_{\text{W}}$$\times 10^{-2}$, computed over $N_{\text{r}}=25$ repetitions, comparing HF-only GLaM and MF-GLaM models for four HF training set sizes. Baseline $\epsilon_{\text{W}}$, for the LF-only GLaM: $10.8\,(5.3)\times 10^{-2}$.
@@ 5 Discussion and conclusions
Across three increasingly complex numerical examples, comprising two synthetic setups and one realistic earthquake application, the MF-GLaM consistently achieves faster convergence to the true HF conditional distributions compared to single-fidelity GLaM stochastic emulators trained only on HF or LF data. Pointwise comparisons of the conditional PDFs at different inputs show that MF-GLaM captures the bulk and shape of the HF distributions more accurately, especially in the presence of limited HF training data. Moreover, incorporating LF runs can substantially reduce the computational cost of emulating HF stochastic simulators. In the earthquake application, combining a modest HF training set with $1,\!000$ LF runs achieved performance comparable to training solely with a significantly larger HF dataset, resulting in approximately 28% total cost savings. In practical scenarios, the HF model can be significantly more computationally expensive than the LF model, often by multiple orders of magnitude (as opposed to the $5.7:1$ ratio in our example), which can make the efficiency gains offered by our framework even more substantial. The framework is equally applicable when the LF and HF responses are defined on different scales, e.g., when the LF output is a correlated but distinct quantity of interest. In such cases, the data should be rescaled to a common scale before fitting the MF-GLaM, since the parameter-level fusion in MF-GLaM does not explicitly account for scale mismatches.
@@ Appendix A Maximum likelihood estimator for the MF-GLaM
We aim to minimize the Kullback–Leibler (KL) divergence between the joint distributions $f_{0}$, given by Eq. 27, and $f_{\mathrm{GLaM}}$, given by Eq. 28. This KL divergence is expressed as:
@@
Although we could estimate the parameters directly from Eq. 43 by replacing the two expectations with empirical means from the HF and LF samples, we instead aim to derive a log-likelihood estimator for $\boldsymbol{\theta}$ in a way to to associate each sample an additional weight, in order to ensure consistency with the Bayesian derivation of the BIC, which we will present later. The weights are obtained via importance sampling. When generating samples within a multifidelity setting, we would typically not sample $S$ following the Bernoulli distribution with $p$ but with a much higher probability $p^{\prime}$ towards the low-fidelity model. For a design of $N_{\text{L}}$ model evaluations for the LF model and $N_{\text{H}}$ for the HF model, the sampling probability can be regarded as $p^{\prime}=\frac{N_{\text{L}}}{N_{\text{L}}+N_{\text{H}}}$. Therefore, we consider an instrumental distribution
@@ Appendix B Bayesian information criterion for MF-GLaMs
Let us denote as $\boldsymbol{D}$ the random vector $(S,\boldsymbol{X},Y_{\text{H}},Y_{\text{L}})$ and $\boldsymbol{{\mathcal{D}}}=({\mathcal{S}},\boldsymbol{{\mathcal{X}}}_{\text{H}},\boldsymbol{{\mathcal{X}}}_{\text{L}},{\mathcal{Y}}_{\text{H}},{\mathcal{Y}}_{\text{L}})$ the entire collection of the observed HF and LF data. As a reminder, we denote the full set of parameters to be estimated as $\boldsymbol{\theta}=(\boldsymbol{c},\boldsymbol{d})$. As seen from Eq. 30, the parameters $\boldsymbol{c},\boldsymbol{d}$ are estimated jointly from the HF and LF data, with $\boldsymbol{c}$ depending on both HF and LF data, while $\boldsymbol{d}$ depending only on the HF data.
@@
where $f_{\boldsymbol{D}}(\boldsymbol{{\mathcal{D}}})$ denotes the marginal distribution of $\boldsymbol{{\mathcal{D}}}$ and $L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})$ is the likelihood of the parameters for the model ${\mathcal{M}}_{k}$ given the data, such that $\log{L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})}=\ell(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})$ in Eq. 30:
@@
which arises from Eq. 50 with $p=0.5$, as discussed in Section 3.2. Then, the posterior probability for model ${\mathcal{M}}_{k}$ is
[…]
To select the model that best describes the observed data, we aim at maximizing this probability with respect to $k$, or equivalently, minimizing $-\log\mathbb{P}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})$. The latter reads:
[…]
@@
[…]
To approximate the integrand in Eq. 57, we expand the log-likelihood about the vector of estimates ${\boldsymbol{\theta}}^{*}_{k}$ obtained by maximizing $\log L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})$. Using second-order Taylor series expansion, we have:
@@
If we substitute the log-likelihood from Eq. 30 into Eq. 61, the Hessian matrix reads:
@@
Substituting $\mathbf{H}$ back into the log-likelihood expansion in Eq. 60 gives:
@@
Then, the information criterion in Eq. 57 becomes:
[…]
@@
[…]
We notice that the integrand of the integral on the right side of Eq. 73 can be directly compared with the multivariate Gaussian PDF. Thus, the integral equals:
[…]
@@
Substituting this back into the right-hand side of Eq. 73 gives:
@@
The right-hand side of Eq. 77 is the BIC estimate for an MF-GLaM ${\mathcal{M}}_{k}$, which can equivalently be written as:
[…]

### v2

@@ 4.3 Multi-story building subject to an earthquake
In our last application, we aim to explore the applicability and performance of our framework in a real-world application involving real earthquake simulations, adopted from Zhu et al. (2023); Schär et al. (2025). More precisely, we consider a three‐story steel frame structure subjected to stochastic seismic excitations. The structure is modeled as a three-degree-of-freedom system and is represented by a computational model, and the input is given by synthetic ground motion time series generated from a stochastic ground motion model fitted to a dataset of real ground motions. In this framework, the quantity of interest (QoI) is the maximum interstory drift, defined as the maximum of the individual interstory drifts, $\Delta_{1}(t),\Delta_{2}(t),\Delta_{3}(t)$, computed for each story during the seismic event:
@@
For the synthetic ground motion model, we adopt the model from Zhu et al. (2023), based on the models from Broccardo and Dabaghi (2017); Rezaeian and Der Kiureghian (2010). The uncertain seismic input is characterized by four ground motion parameters: the expected Arias intensity $I_{a}$, the time $t_{\text{mid}}$ at which 45% of the expected Arias intensity is reached, the effective duration $D_{5-95}$ of the ground motion, and the main frequency $\omega_{g}$ of the motion at $t_{\text{mid}}$. The ground motion parameters are statistically modeled using log-normal marginal distributions coupled with a Gaussian copula, with the details of the distributions shown in Table˜6. As the four ground‐motion parameters are high‐level descriptors of an earthquake signal, they do not uniquely define a full time‐series ground motion. Thus, a single parameter set can lead to significantly different possible signals.
@@
$I_{a}$ [m/s] $\mathcal{LN}\bigl($-4.61, $2.0920\bigr)$ $\begin{pmatrix}1&0.015&$-0.23$&$-0.13$\\ 0.015&1&0.68&$-0.36$\\ $-0.23$&0.68&1&$-0.11$\\ $-0.13$&$-0.36$&$-0.11$&1\end{pmatrix}$
$t_{\mathrm{mid}}\ [\text{s}]$ $\mathcal{LN}\bigl(2.55,\;0.7998\bigr)$
$D_{5\text{--}95}\ [\text{s}]$ $\mathcal{LN}\bigl(2.67,\;0.2826\bigr)$
$\omega_{g}\ [\text{rad/s}]$ $\mathcal{LN}\bigl(1.42,\;0.3475\bigr)$
@@
In our study, different levels of fidelity are regulated by the time step of the simulations. The HF simulator uses a fine time step of $\delta t_{\text{H}}=0.01\penalty 10000\ \text{s}$, whereas the LF simulator is derived from the same model by increasing, but with a seven times larger time step $\delta t_{\text{L}}=0.07\penalty 10000\ \text{s}$. The average wall‑clock runtime for an HF simulation is $\bar{t}_{\text{H}}=10.3\,\penalty 10000\ \text{s}$, whereas the average runtime for an LF simulation is $\bar{t}_{\text{L}}=1.8\,\text{s}$, meaning that an LF simulation is about 5.7 times faster than an HF simulation.
@@
As mentioned above, the four input parameters are statistically dependent and coupled via a Gaussian copula (see Table˜6). However, for the PCEs used in constructing all GLaMs, the inputs are assumed to be independent, as detailed in Section˜2. In such cases, it is common practice to disregard the input dependencies during model construction, as this often leads to more accurate pointwise predictions in data-driven settings compared to transforming the correlated inputs into uncorrelated ones via the highly nonlinear Rosenblatt transform (Torre et al., 2019). However, the input dependencies are still taken into account when generating training and validation samples.
Fig.˜7 compares the empirical response distributions from our true HF and LF simulators to the corresponding predicted distributions from the HF‐only GLaM (red curve) and the MF‐GLaM (blue curve). Three random input samples obtained using LHS are shown in each row, and each column corresponds to a different HF training set size. For each input sample, we generated $N_{\text{R}}=500$ HF runs and $N_{\text{R}}=500$ LF runs to estimate the true HF (black dashed curve) and LF (gray dotted curve) distributions, respectively. The output has been multiplied by 1,0001,\!000 for convenient display in millimeters.
Despite the LF distribution often diverging from the corresponding HF response, the MF‐GLaM successfully captures the bulk and shape of the HF distribution across various scenarios, especially when HF data are scarce. In contrast, the HF‐only GLaM tends to be more sensitive to limited HF samples and can misalign with the empirical HF distribution, e.g., at $\boldsymbol{x}_{2}$ for $N_{\text{H}}=100$ or at $\boldsymbol{x}_{3}$ for $N_{\text{H}}=400$.
@@
Fig.˜8 and the corresponding summary in Table˜7 present the normalized Wasserstein distance $\epsilon_{W}$ for three surrogate models, LF-only, HF-only, and MF-GLaM, across increasing high-fidelity training set sizes from $N_{\text{H}}=100$ to $N_{\text{H}}=800$. The reference validation data were generated by running our HF simulator on 200 Latin hypercube samples, each with $N_{\text{R}}=250$ replications, resulting in a total of 50,00050,\!000 HF runs. In this application, the LF simulator is a numerically less refined version of the HF model, rather than a physically distinct one. This may explain why the LF-only GLaM performs relatively well at small HF training sizes ($N_{\text{H}}\leq 200$), when the available HF data are too limited to fully capture the underlying variability. Overall, the MF-GLaM (blue boxes) consistently achieves the lowest error among the three emulators, with especially large performance gains at smaller HF sample sizes ($N_{\text{H}}\leq 400$).
Accounting for the computational cost of HF and LF runs, where 5.7 LF runs are considered equivalent in cost to one HF run, the total cost of using $N_{H}=200$ HF samples and $N_{L}=1,\!000$ LF samples corresponds to $200+\frac{1,\!000}{5.7}\approx 375$ HF-equivalent runs. This configuration leads to an MF-GLaM with median $\epsilon_{\text{W}}=0.075$ (see Table˜7). This error is significantly lower than the median $\epsilon_{\text{W}}$ of the HF-only GLaM trained on $N_{\text{H}}=400$ HF points, for which $\epsilon_{\text{W}}=0.094$, indicating that the MF approach performs better at a reduced total cost. A similar trend is observed for the MF-GLaM trained on $N_{H}=400$ and $N_{L}=1,\!000$ points, where the total cost amounts to $400+\frac{1,\!000}{5.7}\approx 575$ HF-equivalent runs. In this case, the MF-GLaM’s median error, $\epsilon_{\text{W}}=0.056$, is comparable to that of the HF-only GLaM trained on $N_{\text{H}}=800$ HF points, for which $\epsilon_{\text{W}}=0.059$. Thus, comparable performance is achieved at a significantly lower computational cost, corresponding to a budget saving of approximately $28\%$.
Figure 8: Multi-story building subject to an earthquake. Comparison of the normalized Wasserstein distance $\epsilon_{\text{W}}$ for LF-only, HF-only, and MF GLaMs across increasing HF experimental design sizes $N_{\text{H}}$. The box plots correspond to 25 repetitions of the procedure, and $\epsilon_{\text{W}}$ is computed using 200200 test points, with 250250 replications performed at each point.
Table 7: Multi-story building subject to an earthquake. Median (interquartile range) of the normalized Wasserstein distance $\epsilon_{\text{W}}$ $\times 10^{-2}$, computed over $N_{\text{r}}=25$ repetitions, comparing HF‑only GLaM and MF‑GLaM models for four HF training set sizes. Baseline $\epsilon_{\text{W}}$, for the LF‑only GLaM: $10.8\,(5.3)\times 10^{-2}$.
@@ 5 Discussion and conclusions
Across three increasingly complex numerical examples, comprising two synthetic setups and one realistic earthquake application, the MF-GLaM consistently achieves faster convergence to the true HF conditional distributions compared to single-fidelity GLaM stochastic emulators trained only on HF or LF data. Pointwise comparisons of the conditional PDFs at different inputs show that MF-GLaM captures the bulk and shape of the HF distributions more accurately, especially in the presence of limited HF training data. Moreover, incorporating LF runs can substantially reduce the computational cost of emulating HF stochastic simulators. In the earthquake application, combining a modest HF training set with 1,0001,\!000 LF runs achieved performance comparable to training solely with a significantly larger HF dataset, resulting in approximately 28% total cost savings. In practical scenarios, the HF model can be significantly more computationally expensive than the LF model, often by multiple orders of magnitude (as opposed to the $5.7:1$ ratio in our example), which can make the efficiency gains offered by our framework even more substantial. The framework is equally applicable when the LF and HF responses are defined on different scales, e.g., when the LF output is a correlated but distinct quantity of interest. In such cases, the data should be rescaled to a common scale before fitting the MF-GLaM, since the parameter-level fusion in MF-GLaM does not explicitly account for scale mismatches.
@@ Appendix A Maximum likelihood estimator for the MF-GLaM
We aim to minimize the Kullback–Leibler (KL) divergence between the joint distributions $f_{0}$, given by Eq.˜27, and $f_{\mathrm{GLaM}}$, given by Eq.˜28. This KL divergence is expressed as:
@@
Although we could estimate the parameters directly from Eq.˜43 by replacing the two expectations with empirical means from the HF and LF samples, we instead aim to derive a log-likelihood estimator for $\boldsymbol{\theta}$ in a way to to associate each sample an additional weight, in order to ensure consistency with the Bayesian derivation of the BIC, which we will present later. The weights are obtained via importance sampling. When generating samples within a multifidelity setting, we would typically not sample $S$ following the Bernoulli distribution with $p$ but with a much higher probability $p^{\prime}$ towards the low-fidelity model. For a design of $N_{\text{L}}$ model evaluations for the LF model and $N_{\text{H}}$ for the HF model, the sampling probability can be regarded as $p^{\prime}=\frac{N_{\text{L}}}{N_{\text{L}}+N_{\text{H}}}$. Therefore, we consider an instrumental distribution
@@ Appendix B Bayesian information criterion for MF-GLaMs
Let us denote as $\boldsymbol{D}$ the random vector $(S,\boldsymbol{X},Y_{\text{H}},Y_{\text{L}})$ and $\boldsymbol{{\mathcal{D}}}=({\mathcal{S}},\boldsymbol{{\mathcal{X}}}_{\text{H}},\boldsymbol{{\mathcal{X}}}_{\text{L}},{\mathcal{Y}}_{\text{H}},{\mathcal{Y}}_{\text{L}})$ the entire collection of the observed HF and LF data. As a reminder, we denote the full set of parameters to be estimated as $\boldsymbol{\theta}=(\boldsymbol{c},\boldsymbol{d})$. As seen from Eq.˜30, the parameters $\boldsymbol{c},\boldsymbol{d}$ are estimated jointly from the HF and LF data, with $\boldsymbol{c}$ depending on both HF and LF data, while $\boldsymbol{d}$ depending only on the HF data.
@@
where $f_{\boldsymbol{D}}(\boldsymbol{{\mathcal{D}}})$ denotes the marginal distribution of $\boldsymbol{{\mathcal{D}}}$ and $L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})$ is the likelihood of the parameters for the model ${\mathcal{M}}_{k}$ given the data, such that $\log{L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})}=\ell(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})$ in Eq.˜30:
@@
which arises from Eq.˜50 with $p=0.5$, as discussed in Section˜3.2. Then, the posterior probability for model ${\mathcal{M}}_{k}$ is
$\mathbb{P}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})=f_{\boldsymbol{D}}(\boldsymbol{{\mathcal{D}}})^{-1}\pi({\mathcal{M}}_{k})\int L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})g(\boldsymbol{\theta}_{k}\mid{\mathcal{M}}_{k})\,d\boldsymbol{\theta}_{k}.$ (55)
To select the model that best describes the observed data, we aim at maximizing this probability with respect to $k$, or equivalently, minimizing $\log\mathbb{P}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})$. The latter reads:
$\log\mathbb{P}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})=\log f_{\boldsymbol{D}}(\boldsymbol{{\mathcal{D}}})-\log\pi({\mathcal{M}}_{k})-\log\left\{\int L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})g(\boldsymbol{\theta}_{k}\mid{\mathcal{M}}_{k})\,d\boldsymbol{\theta}_{k}\right\}.$ (56)
@@
$\text{IC}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})=-\log\pi({\mathcal{M}}_{k})-\log\left\{\int L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})g(\boldsymbol{\theta}_{k}\mid{\mathcal{M}}_{k})\,d\boldsymbol{\theta}_{k}\right\}.$ (57)
To approximate the integrand in Eq.˜57, we expand the log-likelihood about the vector of estimates ${\boldsymbol{\theta}}^{*}_{k}$ obtained by maximizing $\log L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})$. Using second-order Taylor series expansion, we have:
@@
If we substitute the log-likelihood from Eq.˜30 into Eq.˜61, the Hessian matrix reads:
@@
Substituting $\mathbf{H}$ back into the log-likelihood expansion in Eq.˜60 gives:
@@
Then, the information criterion in Eq.˜57 becomes:
$\text{IC}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})=-\log\pi({\mathcal{M}}_{k})-\log\left\{\int L(\boldsymbol{\theta}_{k}\mid\boldsymbol{{\mathcal{D}}})g(\boldsymbol{\theta}_{k}\mid{\mathcal{M}}_{k})\,d\boldsymbol{\theta}_{k}\right\}$
$\approx-\log\pi({\mathcal{M}}_{k})-L(\boldsymbol{\theta}^{*}_{k}\mid\boldsymbol{{\mathcal{D}}})\int\exp\left\{-\tfrac{1}{2}(\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}_{k})^{\top}\left[\frac{N_{\text{H}}+N_{\text{L}}}{2}\bar{I}\right](\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}_{k})\right\}g(\boldsymbol{\theta}_{k}\mid{\mathcal{M}}_{k})\,d\boldsymbol{\theta}_{k}$ (72)
@@
$\text{IC}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})\approx-\log\pi({\mathcal{M}}_{k})-L(\boldsymbol{\theta}^{*}_{k}\mid\boldsymbol{{\mathcal{D}}})\int\exp\left\{-\tfrac{1}{2}(\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}_{k})^{\top}\left[\frac{N_{\text{H}}+N_{\text{L}}}{2}\bar{I}\right](\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}_{k})\right\}\,d\boldsymbol{\theta}_{k}$ (73)
We notice that the integrand of the integral on the right side of Eq.˜73 can be directly compared with the multivariate Gaussian PDF. Thus, the integral equals:
$\int\exp\left\{-\tfrac{1}{2}(\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}_{k})^{\top}\left[\frac{N_{\text{H}}+N_{\text{L}}}{2}\bar{I}\right](\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}_{k})\right\}\,d\boldsymbol{\theta}_{k}$ $=(2\pi)^{\frac{n_{\boldsymbol{\theta}_{k}}}{2}}\left|\frac{N_{\text{H}}+N_{\text{L}}}{2}\bar{I}\right|^{-\tfrac{1}{2}}$
@@
Substituting this back into the right-hand side of Eq.˜73 gives:
@@
The right-hand side of Eq.˜77 is the BIC estimate for an MF-GLaM ${\mathcal{M}}_{k}$, which can equivalently be written as:
$\text{MF-BIC}({\mathcal{M}}_{k}\mid\boldsymbol{{\mathcal{D}}})=-2\log L(\boldsymbol{\theta}^{*}_{k}\mid\boldsymbol{{\mathcal{D}}})+n_{\boldsymbol{\theta}_{k}}\log\left(\frac{N_{\text{H}}+N_{\text{L}}}{2}\right).$ (78)